internal energy

One mole of H2O(g) at 1.00 atm and 100 degrees Celsius occupies a volume of 30.6 L. When one mole of H2O(g) is condensed to one mole of H2O(l) at 1.00 atm and 100 degrees Celsius, 40.66 kJ of heat is released. If the density of H2O(l) at this temperature and pressure is 0.996 g/cm3, calculate ΔE for the condensation of one mole of water at 1.00 atm and 100 degrees Celsius.
Answer #1

Given Data:

  • The initial pressure of the water gas is 1.00 atm.

  • The volume of the water gas is 30.6 L.

  • The temperature is 100 degrees Celsius.

  • The heat released is 40.66 kJ.

  • The density of the liquid water is 0.996 g/centimetre cube.


The internal energy is determined by the following formula:

ΔU=q+w...(I)

Here,

  • U is the internal energy.

  • q is the heat.

  • w is the work.


Use the density formula to determine the volume of liquid water as follows:

The mass of one mole water is 18 g/mol.

Density=MassVolume

Substituted the known values.

0.996g/cm3=18gVolumeVolume=18g0.996g/cm3=18.07cm3


Convert the volume of liquid water from centimetre cube to liter as follows:

1000cm3=1L18.07cm3=0.01807L


The work is determined as follows:

w=P(Vliq.waterVgas.water)

Substitute the known values.

w=1.00atm(0.01807L30.6L)=30.58atmL


Convert the work form atm L to joule as follows:

0.0821atmL=8.314J30.58atmL=8.314J0.0821atmL×30.58atmL=3096.73J


Convert the work form joule to kilojoule as follows:

1000J=1kJ3096.73J=3.09673kJ


Use the equation (I) to calculate the internal energy change.

ΔU=40.66kJ+3.09673kJ=37.56kJ


Therefore, the internal energy change for the condensation is -37.56 kJ.


answered by: Ravi Prabhat